In a recent poll of 1200 randomly selected adult office workers, 32% said they had worn a Halloween costume to the office at least once. 1) What is the margin of error, using a 95% confidence level, for estimating the true population proportion of adult office workers who have worn a Halloween costume to the office at least once? (Round to the nearest thousandth) 1) A) 0.026 B) 0.158 C) 0.013 D) 0.053
Answer #1
sample success x = 384
sample size          n= 1200
sample proportion p̂ =x/n= 0.3200
std error se= √(p*(1-p)/n) = 0.0135
for 95 % CI value of z= 1.960
margin of error E=z*std error   = 0.026

option A is correct